# Understanding New-Knowledge-Induced Factual Hallucinations in LLMs: Analysis, Solution and Interpretation

### v1

@@ 4.2 Knowledge-based Reasoning
We investigate how training on a knowledge-based reasoning task with unknown knowledge affects performance across different downstream tasks. Specifically, we examine six distinct test groups: (1) Same-Type Same-Reasoning (STSR): the exact reasoning task that trained with unknown knowledge type; (2) Same-Type Different-Reasoning (STDR): different reasoning tasks within the same knowledge type; (3) Different-Type Different-Reasoning (DTDR): all other reasoning tasks with different knowledge; (4) Same-Type QA (STQA): the QA task with the same knowledge type as STSR; (5) Different-Type QA (DTQA): QA tasks with other three knowledge types; and (6) wiki: the real-world QA test set.
@@
We measure the relative performance change with respect to the baseline and compute the average difference within each of the six task groups. Results in Figure 3 show that learning reasoning tasks with new knowledge consistently induces performance degradation across all six groups. The overall trend aligns with previous findings: the most severe hallucinations occur in STSR, indicating strong intra-task interference, while hallucination on other task groups are relatively minor. A notable difference is that among other tasks, QA test sets exhibit even stronger hallucinations than seemingly more related reasoning tasks: STQA, DTQA, and even wiki show greater degradation than STDR and DTDR. Within QA, the degradation is more pronounced in STQA than in DTQA, implying that shared factual grounding heightens vulnerability.
@@
Figure 3: The impact of learning new knowledge in reasoning tasks on the average performance across different groups. Results of each variant model on each dataset are presented in Appendix C.
5 KnownPatch: Mitigating Hallucination
In Section 4.1, we find that when all factual knowledge in a certain type is new, even if it constitutes a small proportion of the overall data, it significantly increases the risk of hallucination.
Prior work typically attempts to mitigate hallucinations by filtering out all data that contain unknown knowledge. However, in real-world settings, factual knowledge spans multiple categories and is intertwined with diverse task contexts, making it difficult to separate knowledge from task-related information. Identifying which portions of the data involve unknown knowledge is often ambiguous, and discarding entire samples that might contain it can be highly inefficient. As a result, perfect filtering is neither feasible nor effective in practice, and the remaining new knowledge can still induce serious hallucinations.
Even if perfect filtering were achievable, one might expect that transferring the new knowledge to Continued Pre-Training (CPT) would allow the model to first acquire it safely before SFT. However, our experiments show that learning new knowledge during the CPT phase can also cause hallucination behaviors similar to the observations in Section 4.1, with detailed results provided in Appendix E.
To address this issue, we propose KnownPatch, a simple yet effective approach that stabilizes the model by briefly reinforcing known knowledge at the final stage of training.
5.1 Intuition
We observe that learning tasks with known knowledge does not introduce instability, whereas training on unknown knowledge can easily disrupt the model’s behavior. Based on this observation, KnownPatch places a small portion of known knowledge at the very end of the training sequence. The idea is to allow the model, after being inevitably disturbed by new knowledge, to recover a stable state through a brief reinforcement of previously known information.
This approach is lightweight and easy to apply, as it only requires identifying a small subset of known samples from the SFT corpus. While precise separation of all known and unknown knowledge is practically infeasible, extracting a small subset of clearly known data is straightforward (Gekhman et al. 2024), making KnownPatch simple yet effective in stabilizing the overall training process.
5.2 Setups
KnownPatch modifies the training sequence by placing a small subset of SFT data with all known knowledge at the end of the training. In our setting, the data trained prior to this subset consists entirely of unknown knowledge across all of the types, to simulate the worst situation.
We consider two scenarios: (1) the ideal scenario, where the injected known knowledge covers all knowledge types that are involved during the previous training; (2) a more realistic scenario, where the injected known knowledge does not cover all types, and we specifically examine the case where known data from only one type is missing.
We evaluate KnownPatch across injection ratios of 5%, 10%, and 20% under both scenarios. The baseline model is trained on randomly shuffled data, where most knowledge within each type remains unknown; the theoretical upper bound is a model trained solely on known knowledge with the same dataset size.
To present the observations more clearly in the main text, we average and aggregate certain results, while detailed results are provided in Appendix F. In the first scenario, we average the results across the four test types (B, D, M, U) and collectively refer to them as QA, while the OOD wiki results, which are not involved in training, are reported separately. In the second scenario, we train four variant models, each corresponding to a different missing type in the injected known knowledge. We then average the performance of each variant on its respective missing-type test set and refer to this average as QA, while the averaged wiki performance across the four variants is reported separately.
5.3 Results
Figure 4: Averaged performance and attention score changes under different settings. The two bar groups on the left represent the performance of KnownPatch in the first scenario, while the bar group on the right corresponds to the second scenario. Values here represent the performance change percentage compared to the upper-bound model.
KnownPatch effectively mitigates factual hallucinations with minimal data injection. The left two groups of bars in Figure 4 show the performance of KnownPatch in knowledge QA tasks when injecting 5% and 20% of known data at the end of training. When the injected data covers all previous unknown knowledge types, KnownPatch consistently outperforms the baseline with randomly shuffled data across all injection ratios. Even with only 5% injected known data, KnownPatch recovers a substantial portion of the performance lost due to hallucinations. At 20%, the model’s performance becomes very close to that of the upper-bound model trained entirely on known knowledge. Notably, on the OOD wiki test set, performance with just 5% injected known data already approaches the upper bound, and at 20%, it even exceeds the performance of the fully known model, suggesting that late-stage injection of diverse known facts may enhance generalization and stabilize cross-domain knowledge application.
Even when one knowledge type is missing from the injection data, KnownPatch still substantially mitigates hallucination. As shown in the right group of bars in Figure 4, the model achieves significant performance gains even on the uncovered type, despite never being trained on its known facts. On the OOD wiki test set, performance also shows a clear improvement over the shuffled baseline and even comparable to the all-known upper-bound. This suggests that KnownPatch induces a global stabilization effect that benefits even types not directly represented in the patch.
KnownPatch also performs well in reasoning tasks. In knowledge-based reasoning tasks, when the injected patch data covers all previously unknown knowledge-based reasoning tasks, KnownPatch also achieves consistent performance improvements across all task types with different injection ratios. This effect can also transfer to QA tasks that are affected but not directly injected. Figure 5 shows the results of 20%, demonstrating the general effectiveness of the KnownPatch method.
Figure 5: Performance of KnownPatch on reasoning tasks when injecting 20% known data. The values here represent the accuracy percentage of this model compared to the fully known upper-bound model.
6 Interpretability with Attention Analysis
In this section, we analyze the mechanisms under new-knowledge-induced factual hallucinations through the attention scores. We measure the relative changes in entity attention and task performance in models trained with different data compositions (using the model trained with entirely known data as the baseline). We also analyze the attention pattern changes after applying the KnownPatch method.
6.1 Analysis Setup
@@
We analyze the changes in attention scores that the model assigns to key entities in the context when applying knowledge. In the Biography-Reasoning dataset, the key entity in each question is the person name. Therefore, we examine the model’s attention score on the names when generating the first token of the knowledge.
Prior interpretability studies suggest that the model processes inputs in three stages: understanding, execution, and output, among which language-agnostic knowledge retrieval and abstract reasoning primarily occur in the middle-to-later layers (Wendler et al. 2024; Zhao et al. 2024). We also examine the model’s average attention to entities across different layers in both reasoning and knowledge QA tasks (detailed in Appendix D), with results shown in Figure 6, demonstrating that this attention is significantly higher in layers 12–24 (out of 28 total layers in Qwen2.5-1.5B), consistent with prior works. Therefore, in this study, we compute the average attention scores over layers 12 to 24 for interpretability analysis. We measure the relative change in entity attention to assess how training on new knowledge affects the model’s focus on critical entities, by comparing models trained under some new data to the model trained entirely on known knowledge, i.e. the upper bound model.
@@
Figure 6: Attention score on the target name across layers in QA and reasoning training setups. The solid curves show the average attention score at each layer, aggregated across all datasets and instances. The shaded regions represent the standard deviation.
We use this method to analyze (1) the performance degradation of QA tasks under different proportions of unknown knowledge (Figure 2), (2) the impact of learning new knowledge in reasoning tasks (Figure 3), and (3) the effect of applying KnownPatch (Figure 4).
6.2 Analysis Results and Observations
Figures 7, 8 and 9 show the changes in entity attention and task performance for models trained on unknown knowledge in various settings.
@@ 6.2 Analysis Results and Observations
Hallucinations emerge alongside a concurrent decline in entity attention and performance, and this coupling effect intensifies in the absence of known knowledge. Figure 7 presents the interpretability analysis corresponding to Figure 2. As the proportion of unknown instances within a knowledge type increases, the model’s attention to key entities gradually declines, accompanied by more severe hallucinations. The decline in entity attention under RemoveKnown is faster and leads to a more abrupt performance drop than under KeepKnown, indicating that the absence of known information within a knowledge type accelerates attention decay, which is strongly correlated with performance degradation.
@@
Contextual similarity plays a crucial role in how hallucinations propagate across tasks. Figure 8 presents the interpretability analysis corresponding to Figure 3. We observe that both the performance drop and the reduction in attention to key entities are most pronounced in STSR, followed by the QA tasks. Overall, these two metrics exhibit a strong correlation across different task groups. Table 3 further shows the contextual similarity between STSR and other task groups. On average, QA tasks demonstrate the highest contextual similarity to the reasoning task involving unknown knowledge. Owing to their relatively long input contexts, reasoning tasks tend to be less similar to each other, whereas knowledge QA contexts often overlap with substrings of reasoning contexts. Taken together, these findings indicate that performance degradation grows with contextual similarity, implying that inter-task hallucinations are likely to spread through shared contextual patterns. Consequently, learning new knowledge in reasoning tasks is less likely to induce hallucinations in other reasoning tasks, while QA tasks are more vulnerable to such interference.
Figure 8: Accuracy and attention score changes compared to the all-known baseline model when learning new knowledge in reasoning tasks.
@@
1.0000 0.6164 0.5896 0.7312 0.6982 0.7199
Table 3: Averaged contextual similarity between STSR and other test groups, where each value denotes the mean similarity between the STSR context and all task contexts within the corresponding group. Similarity between A and B is computed as the proportion of tokens in B appearing in A.
KnownPatch stabilizes the model’s attention patterns and mitigates hallucinations. Figure 9 shows the performance drop and changes in attention on QA test sets relative to the all-known baseline model, providing an interpretability analysis corresponding to the right group of bars in Figure 4. It can be observed that KnownPatch can effectively restore the model’s focus on key entities in the questions, thereby significantly alleviating the hallucination effects caused by learning new knowledge.
Figure 9: Performance and attention score changes when learning new knowledge in QA tasks, and after applying KnownPatch (with 20% known data). QA represents the average across the four QA test sets, and error bars indicate standard deviations.
@@ 7 Conclusion
In this work, we present a systematic study on hallucinations caused by learning new knowledge in LLMs, examining their behavior across knowledge types and task types. Our experiments reveal that even a small number of fully unknown facts can trigger severe hallucinations, not only in the target task but also in other unrelated knowledge QA tasks, indicating a non-trivial inter-task impact. To address this hallucination, we propose KnownPatch, a simple but practical mitigation strategy that injects a small proportion of known knowledge samples during the final phase of fine-tuning. Without extensive data, KnownPatch effectively reduces hallucination across both QA and reasoning tasks even under partial type coverage. Through further analysis, we find that learning new knowledge shifts the model’s focus from key entities to other context, promoting erroneous binding and hallucination propagation. And this altered attention pattern can propagate along similar contexts. KnownPatch counteracts this by restoring entity-centric attention patterns.
@@ A.1 Wiki Details
The ENTITYQUESTIONS dataset from Wikidata is divided into multiple subsets, such as P17, P20, etc., with each subset containing questions of the same format. For example, an instance from P17 is “Which country is Juniper Bank located in?”, and an instance from P20 is “Where did Connee Boswell die?”. Based on Gekhman et al. (2024)’s classification of knowledge, we categorize Wikipedia knowledge into four levels: HighlyKnown, MaybeKnown, WeaklyKnown, and Unknown. We construct the test set using subsets of HighlyKnown and MaybeKnown instances. To ensure the balanced distribution of the test set, we sample approximately the same number of questions from each subset, resulting in a final test set of 1,000 questions.
@@ A.2 Biography-Reasoning Details
The first name and last name of each individual are selected from separate pools and are ensured to be unique. For first names, we use 3,000 English names with an equal split between male and female names (this affects the use of gendered pronouns in reasoning tasks). For last names, we select 250 Chinese surnames , which are then randomly paired with the first names in a balanced manner. This random combination of English first names and Chinese last names is designed to generate synthetic individuals that minimize overlap with real-world knowledge already known to language models.
@@
To construct knowledge reasoning tasks, we introduce a set of auxiliary knowledge. Specifically, our dataset involves relations such as major → field (e.g., Computer Science belongs to Engineering) and university → country (e.g., Stanford University belongs to the United States). These auxiliary facts already exist in the model’s pre-trained knowledge base. To ensure that the model reliably retains them, we also rephrase each auxiliary fact 50 times and include them in the CPT data. All auxiliary knowledge is provided in Tables 5 and 6.
@@
For each attribute of a synthetic individual, we construct three types of reasoning questions. In Table 4, we provide an example for each category of QA and reasoning questions in the dataset.
@@
During the SFT stage for reasoning tasks, we split the known individual set into two subsets: 80% are used to generate reasoning questions, and the remaining 20% are used for QA tasks. This allows the model to be exposed to both task types, ensuring a more reliable evaluation across reasoning and QA.
@@
Table 4: Examples of each QA and reasoning tasks in Biography-Reasoning.

### v2

@@ 4.2 Knowledge-based Reasoning
We investigate how training on a knowledge-based reasoning task with unknown knowledge affects performance across different downstream tasks, including both knowledge-based reasoning and knowledge QA. Specifically, we examine three groups of reasoning tasks: (1) Same-Type Same-Reasoning (STSR): the exact reasoning task that trained with unknown knowledge type; (2) Same-Type Different-Reasoning (STDR): different reasoning tasks within the unknown knowledge type; and (3) Different-Type Different-Reasoning (DTDR): all other reasoning tasks with different knowledge types. We also evaluate the model on the knowledge QA test groups defined in Section 4.1, namely STQA, DTQA, and Wiki.
@@
We measure the relative performance change with respect to the baseline model and compute the average difference within each of the six task groups. Results in Figure 3 show that learning reasoning tasks with new knowledge consistently induces performance degradation across all six groups. The overall trend aligns with previous findings: the most severe hallucinations occur in STSR, indicating strong intra-task interference. Moreover, among other tested tasks, QA test sets exhibit even stronger hallucinations than several seemingly more related reasoning tasks: STQA, DTQA, and even the Wiki test set show greater performance degradation than STDR and DTDR.
@@
Figure 3: The impact of learning new knowledge in reasoning tasks on the average performance across different groups. Detailed results are presented in Appendix C.1.
5 Interpretability Analysis
In this section, we analyze the underlying mechanisms of new-knowledge-induced factual hallucinations by examining the relative changes in attention patterns and hallucination severity. Based on this analysis, we propose a simple training intervention, referred to as KnownPatch, which injects a small amount of known knowledge at the later training stage to restore attention patterns and alleviate hallucination behavior. Crucially, we also investigate the role of contextual similarity, providing evidence that hallucinations propagate primarily through shared lexical contexts.
5.1 Attention Analysis Setup
@@
We measure how learning new knowledge alters the model’s attention to key entities. In the Biography-Reasoning dataset, the key entities are person names, so we quantify the model’s attention to the name tokens when generating the first token of the related knowledge.
Prior interpretability works suggest that knowledge retrieval and reasoning occur primarily in mid-to-late transformer layers (Wendler et al., 2024; Zhao et al., 2024). We confirm this trend in our model by examining attention across layers in both QA and reasoning settings (as detailed in Appendix D). Figure 4 shows that attention on key entities peaks in layers 12-24 (out of 28 layers), so we average attention over these layers in all subsequent analyses. We measure the relative change in entity attention by comparing models trained under new knowledge to the model trained entirely on known knowledge, i.e. the baseline model.
@@
Figure 4: Attention score on the key entity name across layers in QA and reasoning training setups. The solid curves show the average attention score aggregated across all instances. The shaded regions represent the standard deviation.
5.2 Correlation between Attention and Hallucination
STQA DTQA Wiki
Acc. Attn. Acc. Attn. Acc. Attn.
w/o attnKL -54.26 ($\pm$ 3.30) -28.02 ($\pm$ 10.08) -2.67 ($\pm$ 1.58) -5.29 ($\pm$ 4.08) -17.72 ($\pm$ 6.56) -7.33 ($\pm$ 3.03)
w/ attnKL -22.22 ($\pm$ 11.70) -14.18 ($\pm$ 6.09) -5.45 ($\pm$ 3.68) -11.69 ($\pm$ 7.88) -7.13 ($\pm$ 6.57) -10.14 ($\pm$ 1.12)
Table 3: Results of adding KL loss. “Acc.” and “Attn.” denote the averaged performance drop and attention drop on key entities (%, mean with standard deviation), respectively. The baseline model is the same as Section 4.1.
@@ 5.2 Correlation between Attention and Hallucination
Hallucinations correlate with declines in entity attention. Figure 5 presents the interpretability analysis corresponding to Figure 2. As the proportion of unknown instances within a knowledge type increases, the model’s attention to key entities gradually declines, accompanied by more severe hallucinations. Compared to KeepKnown, RemoveKnown exhibits a sharper attention decline and performance drop, indicating that the absence of known information accelerates attention decay and exacerbates performance degradation. Figure 6 shows the interpretability analysis for the reasoning tasks corresponding to Figure 3, where attention patterns are also correlated with hallucinations.
Similar experiments are also conducted on the real-world ENTITYQUESTIONS dataset. The results are reported in Appendix C.2.
@@
Figure 6: Accuracy and attention score changes when learning new knowledge in reasoning tasks.
[…]
Constraining attention alleviates factual hallucinations. Table 3 presents a comparison with Table 2 after introducing the KL constraint. We observe that, on the STQA test sets whose knowledge types are consistent with the unknown knowledge in the training data, as well as on the OOD Wiki test set, hallucinations are substantially alleviated. However, we also find that for DTQA and the wiki test set, attention on key entities dropped more even after applying the KL constraint. To further investigate this phenomenon, we apply the same KL constraint to train another baseline model (trained entirely on known knowledge). This results in an average accuracy drop of 3.67% and a 9.21% reduction in attention to key entities on the test sets, compared to the normally trained baseline model. This indicates that training on known knowledge naturally encourages increased attention to key entities, which we believe is beneficial for learning the task. Consequently, forcing attention to remain unchanged inevitably suppresses this, leading to a modest reduction in entity attention and a corresponding loss in accuracy. Overall, these results suggest that attention shifts during new knowledge training are a key contributing factor to the emergence of hallucinations.
5.3 KnownPatch: Late-stage Injection of Known Knowledge
Building on the above analysis, which shows that training on known knowledge promotes increased attention to key entities, we further explore whether injecting a small amount of known knowledge at the final stage of training can help alleviate factual hallucinations induced by learning new knowledge. We refer to this training method as KnownPatch. The underlying intuition is that exposure to unfamiliar knowledge can disrupt the model’s attention patterns, whereas re-introducing known knowledge encourages the restoration of entity-centric attention and leads to more stable model behavior.
The training data used before applying the fully known injection patch consist entirely of unknown knowledge across all types, simulating a worst-case scenario. We define the injection ratio as the proportion of injected known samples relative to the total training data, and we experiment with injection ratios of 5%, 10%, and 20%. The baseline model is trained on known knowledge of the full training data size, serving as an upper bound. To control for the effect of training order, we additionally compare against a model trained on the same mixed data after shuffling. Our analysis focuses on the relative performance and attention changes of the models trained under the KnownPatch and Shuffled settings compared to the baseline model.
Figure 7: Performance and attention score changes under Shuffled and KnownPatch (with 20% injection ratio) settings. QA represents the average across the four QA test sets, and error bars indicate standard deviations.
KnownPatch stabilizes the model’s attention patterns and mitigates hallucinations. Figure 7 reports the performance drop and the change in entity attention on the QA test sets under a 20% injection ratio. Across both in-domain QA and the OOD wiki set, KnownPatch consistently outperforms the shuffled baseline. With 20% injection, QA performance approaches the all-known upper bound, and performance on the OOD Wiki set even slightly exceeds it. Results under different injection ratios and reasoning tasks are provided in Appendix F, and we find that even with only 5% injection, KnownPatch already yields substantially higher performance than the Shuffled setting on both QA and Wiki test sets. These gains are accompanied by increased attention to key entities in the questions, suggesting that KnownPatch restores entity-centric attention patterns disrupted by training on unknown knowledge and yields a robust mitigation effect that generalizes to OOD data.
5.4 Mechanism of Hallucination Propagation
We further aim to investigate how new-knowledge-induced factual hallucinations propagate across different tasks.
@@ 5.4 Mechanism of Hallucination Propagation
1.00 0.62 0.59 0.73 0.70 0.72
Table 4: Averaged token-level overlap between STSR and other test groups. Each value represents the mean token overlap ratio, defined as the proportion of tokens in a task context that also appear in the STSR context, averaged over all tasks in the corresponding group.
Since attention weights form a normalized distribution that sums to one across all input tokens, a reduction in attention to key entities implies a corresponding shift of attention toward the remaining contextual tokens. Figure 3 shows that learning new knowledge in reasoning tasks has a more substantial impact on QA test sets than on other reasoning test sets. This is consistent with the lexical similarity (measured by token-level overlap) analysis in Table 4: QA test sets are generally more lexically similar to the STSR tasks that trained with unknown knowledge, and this higher lexical similarity aligns with the observed trend of stronger hallucination propagation. This can be attributed to the fact that reasoning tasks typically involve relatively long and diverse input contexts, making them lexically less similar to one another, whereas many knowledge QA contexts share substrings with reasoning trajectories.
Lexical Same Similar Different
Semantic Same Same Diff. Same Diff.
Token Overlap 1.00 0.97 0.89 0.62 0.52
Hallucination 25.57 18.80 16.61 6.65 5.90
Table 5: Averaged contextual similarity (token overlap) and performance drop (hallucination, %).
To disentangle the effects of lexical similarity and semantic similarity in contexts on hallucination propagation, we construct extended variants of the reasoning tasks. Specifically, for each of the 12 reasoning tasks in Biography-Reasoning, we construct four variant tasks under the same knowledge type, defined by lexical similarity (similar vs. different) and semantic similarity (same vs. different). An example for each variant task is provided in Appendix A. We then examine, across the four variant tasks, the extent of performance degradation when the Origin task is trained on unknown knowledge compared to training on known knowledge. Results in Table 5 show that learning new knowledge induces stronger hallucinations on tasks with higher lexical similarity. This confirms that hallucinations propagate primarily through lexical similarity rather than semantic relatedness.
@@ 6 Conclusion
In this work, we present a systematic study on new-knowledge-induced factual hallucinations in LLMs, examining their behavior across knowledge types and task types. Our experiments reveal that even a small number of fully unknown facts can trigger severe hallucinations and can propagate to other tasks. Our analysis reveals that this behavior is closely associated with shifts in attention: learning new knowledge reduces attention to key entities, whereas training on known knowledge reinforces entity-centric attention. Motivated by this observation, we explore a simple training intervention, KnownPatch, which injects a small amount of known knowledge at the late stage of training to restore attention patterns and mitigate hallucinations. Finally, we show that the extent of hallucination propagation increases with lexical similarity between contexts, rather than semantic relatedness, highlighting contextual similarity as a key driver of cross-task hallucination transfer.
Limitations
While our work offers a in-depth analysis of new-knowledge-induced factual hallucinations, there are several boundaries to our current study.
Regarding the synthetic data:
To strictly disentangle new knowledge from facts already internalized during pre-training, we constructed the synthetic Biography-Reasoning dataset. While this controlled environment was necessary, it may not fully capture the noise and semantic complexity of naturally occurring text. We partially mitigated this by validating our findings on the real-world ENTITYQUESTIONS dataset, yet exploring more complex linguistic structures remains a direction for future work.
Regarding model scale:
Our analysis primarily relies on open-source models (e.g., Qwen2.5-1.5B) to enable detailed and fine-grained investigation. While we verify the key findings on larger variants (e.g., Qwen3-8B, Qwen2.5-32B) in the appendix, we acknowledge that extending our experiments to extremely large-scale models (e.g., 70B+ or even larger models) is left for future work due to computational constraints.
@@ A.1 Wiki Details
The ENTITYQUESTIONS dataset from Wikidata is divided into multiple subsets, such as P17, P20, etc., with each subset containing questions of the same format. For example, an instance from P17 is “Which country is Juniper Bank located in?", and an instance from P20 is “Where did Connee Boswell die?". Based on Gekhman et al. (2024)’s classification of knowledge, we categorize Wikipedia knowledge into four levels: HighlyKnown, MaybeKnown, WeaklyKnown, and Unknown. We construct the test set using subsets of HighlyKnown and MaybeKnown instances. To ensure the balanced distribution of the test set, we sample approximately the same number of questions from each subset, resulting in a final test set of 1,000 questions.
@@ A.2 Biography-Reasoning Details
The first name and last name of each individual are selected from separate pools and are ensured to be unique. For first names, we use 3,000 English names from the UCI Machine Learning Repository dataset [^2] https://archive.ics.uci.edu/dataset/591/gender+by+name, which is under CC BY 4.0 license and could be used for any purpose., with an equal split between male and female names (this affects the use of gendered pronouns in reasoning tasks). For last names, we select 250 Chinese surnames from a GitHub repository [^3] https://github.com/smashew/NameDatabases/, which is under ‘The Unlicense’ that allows anyone to use for free. , which are then randomly paired with the first names in a balanced manner. This random combination of English first names and Chinese last names is designed to generate synthetic individuals that minimize overlap with real-world knowledge already known to language models.
@@
We refer to the four attributes Birth, Death, Major and University as B, D, M and U, respectively.
@@
To construct knowledge reasoning tasks, we introduce a set of auxiliary knowledge. Specifically, our dataset involves relations such as major → field (e.g., Computer Science belongs to Engineering) and university → country (e.g., Stanford University belongs to the United States). These auxiliary facts already exist in the model’s pre-trained knowledge base. To ensure that the model reliably retains them, we also rephrase each auxiliary fact 50 times and include them in the CPT data. All auxiliary knowledge is provided in Tables 8 and 9.
@@
For each attribute of a synthetic individual, we construct three types of reasoning questions. We refer to Single Reasoning, Comparative Reasoning, and Novel Reasoning as SR, CR, and NR, respectively. In Table 6, we provide an example for each category of QA and reasoning questions in the dataset.
@@
To investigate the effect of context similarity on hallucination propagation, we construct variant tasks for each reasoning problem under the same knowledge type: (a) Same Meaning, Similar Lexical (SMSL); (b) Same Meaning, Different Lexical (SMDL); and (c) Different Meaning, Similar Lexical (DMSL). Table 7 presents representative examples of these variants. For the Different Meaning, Different Lexical setting, we directly adopt the STDR tasks defined in Section 4.2.
@@
Table 6: Examples of each QA and reasoning tasks in Biography-Reasoning. B, D, M, and U denote birth year, death year, major, and university, respectively; SR, CR, and NR denote Single Reasoning, Comparative Reasoning, and Novel Reasoning, respectively.
Category Example
B_SR SMSL: Is Hakam Cheng’s birth year an odd number?
SMDL: Can Hakam Cheng’s year of birth, when considered as an integer, be classified under the category of odd values?
DMSL: Is the number of Hakam Cheng’s birth year an even number?
B_CR SMSL: How many years is the birth year between Hakam Cheng and Graicyn Xian apart?
SMDL: By what number of years are Hakam Cheng and Graicyn Xian separated in terms of their birth years?
DMSL: How many months apart is the birth year between Hakam Cheng and Graicyn Xian?
B_NR SMSL: What’s the MScore of Hakam Cheng’s birth year?
SMDL: Determine the specific MScore value attributed to the calendar year during which Hakam Cheng was born.
DMSL: What is the AScore of Hakam Cheng’s birth year?
D_SR SMSL: What is the 10th anniversary of Hakam Cheng’s death year?
SMDL: After 10 years from Hakam Cheng passed away, what year does that correspond to?
DMSL: What is the year before the 10th anniversary of Hakam Cheng’s death?
D_CR SMSL: Hakam Cheng or Graicyn Xian, who died first?
SMDL: Identify which individual—Hakam Cheng or Graicyn Xian—died at an earlier date.
DMSL: Who died later, Hakam Cheng or Graicyn Xian?
D_NR SMSL: What’s the AScore of Hakam Cheng’s death year?
SMDL: Which AScore value corresponds to the calendar year in which Hakam Cheng died?
DMSL: What is the MScore of Hakam Cheng’s death year?
M_SR SMSL: Which field does Hakam Cheng’s major belong to?
SMDL: To which academic discipline can the major pursued by Hakam Cheng be categorized?
DMSL: What is the first letter of the field that Hakam Cheng’s major belongs to?
M_CR SMSL: Do Hakam Cheng’s and Graicyn Xian’s majors belong to the same field?
SMDL: Are Hakam Cheng and Graicyn Xian’s areas of academic specialization considered part of the same disciplinary category?
DMSL: Do Hakam Cheng and Graicyn Xian’s majors belong to different field?
M_NR SMSL: What is the sequence of odd-positioned letters within the first word of Hakam Cheng’s major name?
SMDL: From the initial word of Hakam Cheng’s major, extract the characters occupying positions with odd indices and present them in order.
DMSL: What is the sequence of even-positioned letters in the first word of Hakam Cheng’s major name?
U_SR SMSL: Hakam Cheng attend university in which country?
SMDL: Identify the nation within whose borders Hakam Cheng pursued university-level studies.
DMSL: What is the last letter of the country in which did Hakam Cheng attended university?
U_CR SMSL: Hakam Cheng and Graicyn Xian are college alumni?
SMDL: Have Hakam Cheng and Graicyn Xian both completed their higher education at the same university?
DMSL: Are Hakam Cheng and Graicyn Xian not college alumni?
U_NR SMSL: What’s the sequence of the first and last letters of each word in Hakam Cheng’s university name?
SMDL: Provide the ordered sequence formed by the initial and final characters of every word appearing in Hakam Cheng’s university title.
DMSL: What is the sequence of the last and first letters of each word in Hakam Cheng’s university name?
Table 7: Examples of the variant tasks for reasoning tasks.